OpenStudy (trojanpoem):

If z, y were the roots of the equation : x^2 -6 x + 1 = 0 Prove that : (y^n + z^n) is a Positive integer which can be divided by 5 for any integer n

OpenStudy (phi):

I would first solve for the roots. what are they?

OpenStudy (trojanpoem):

3 (+ or -) 2 sqrt(2)

ganeshie8 (ganeshie8):

\[x^2-6x+1=0\implies x^2=6x-1\] Since \(z,y\) are roots, we have \[y^2 = 6y-1\implies y^n=6y^{n-1}-y^{n-2}\tag{1}\] \[z^2=6z-1\implies z^n=6z^{n-1}-z^{n-2}\tag{2}\] Add them both and conclude by induction

OpenStudy (trojanpoem):

Hmm, I have to solve it using Combinations

OpenStudy (trojanpoem):

And permutations

OpenStudy (trojanpoem):

Maybe this can help , The book has this solved example y = x^4 + 6 x^3 + 11 x^2 + 6 x prove that it can be divided by 24 and was solved like this

OpenStudy (trojanpoem):

y = x^4 + x^3 + 5x^3 + 5x^2 + 6 x^2 + 6x y = x^3(x+1) + 5x^2(x+1) + 6x(x+1) y = x(x+1)(x^2 + 5x +6 ) y = x(x+1)(x+2(x+3) = (x+3)P4 (x+3)P4 / 4! = x+3C4 always positive integer so it can be divided by (24 = 4!)

OpenStudy (trojanpoem):

Now, it asks to apply what I learnt in the question above.

ganeshie8 (ganeshie8):

Thats pretty clever!

ganeshie8 (ganeshie8):

Your book is using the fact that \(\binom{n}{r}\) is an integer. You could also view it like this : product of any \(n\) consecutive integers is divisible by \(n!\)

OpenStudy (trojanpoem):

Yeah, I understood from the explanation which is too clear. The problem is to solve "try yourself" problem

OpenStudy (trojanpoem):

Any ideas ?

ganeshie8 (ganeshie8):

Not getting any ideas yet but im still trying to see if the induction method that i suggested before really works

OpenStudy (trojanpoem):

Ok.

OpenStudy (dan815):

binomial theorem

OpenStudy (dan815):

since u are mentioned permutations and combinations

OpenStudy (trojanpoem):

I think you are right.

OpenStudy (dan815):

look at (z+y)^n = ....

OpenStudy (trojanpoem):

Y^n + z^n we will sum them , and maybe the result can help.

OpenStudy (dan815):

what will z^n+y^n be equal to from that binomial expansion

OpenStudy (dan815):

yeah play around with that, i think u can show its divisible by 5 fomr it

OpenStudy (trojanpoem):

I failed , try this before.

OpenStudy (trojanpoem):

I tried*

OpenStudy (trojanpoem):

((3+sqrt(2))^2 + (3-sqrt(2))^2) = 2 Singles

OpenStudy (phi):

** (y^n + z^n) is a Positive integer which can be divided by 5 for any integer n *** I am having a problem with the divide by 5. if we do a simple numerical check with n=2 and y= 3+2 sqr(2), z= 3 - 2 sqr(2) I get y^2 + z^2 = 34

ganeshie8 (ganeshie8):

yeah it doesn even work for n=1

OpenStudy (trojanpoem):

You think 5 is a typo from the book ?

OpenStudy (trojanpoem):

so It's 2 not 5 as 2 will work for any n

OpenStudy (trojanpoem):

Any ideas ?

OpenStudy (dan815):

recheck the question

OpenStudy (phi):

I would assume the problem would be closer to the example you posted, and you start with factors that are consecutive. And then you play the same trick. But that is not the case here, so I have no clear idea of what is the purpose of this question is here.

OpenStudy (trojanpoem):

It's written as I have copied it.

ganeshie8 (ganeshie8):

if psble take a screenshot of as much as possible and attach

OpenStudy (trojanpoem):

The second example was intended to show how the writer of the book thinks and what he wants in answering my problem

OpenStudy (trojanpoem):

I can't transfer my photos to the PC or laptop. But tell me what you mean by as much as possible and I will fetch it.

OpenStudy (dan815):

maybe its x^2 -5 x + 1 = 0 then for every odd n z^n + y^n is divisible by 5

OpenStudy (trojanpoem):

It's written as x^2 -6 x + 1 = 0 in the book. But fine , all I want is to know it's idea solve yours

OpenStudy (dan815):

k lets work on this new problem then x^2 -5 x + 1 = 0 then for every odd positive integer n z^n + y^n is divisible by 5

OpenStudy (dan815):

prove

OpenStudy (trojanpoem):

Lol me ? IDK I am even stuck with 5 more in this book.

OpenStudy (dan815):

oh this is kidn of silly lol, the sqrt 29 in the roots will cancel out for ever odd integer

OpenStudy (dan815):

okay well we need a different question

OpenStudy (trojanpoem):

Post here ?

OpenStudy (dan815):

yeah

OpenStudy (trojanpoem):

Find the coefficient of x^n in the expansion of (1 + x + 2x^2 + 3x^3 + ..... + nx^n)^2

OpenStudy (dan815):

u can write out on liek a dot product and it will be obvious

OpenStudy (trojanpoem):

Show me

OpenStudy (dan815):

uh wait no lemme see

OpenStudy (dan815):